# ISAC Network Planning: Sensing Coverage Analysis and 3-D BS Deployment Optimization

### v1

@@ III-B Scaling Law Analysis
$\frac{({\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime})^{T}}{\|{\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime}\|}=\frac{(\kappa{\bf{b}}_{n}-\kappa\mathbf{t})^{T}}{\kappa\|{\bf{b}}_{n}-\mathbf{t}\|}=\frac{({\bf{b}}_{n}-\mathbf{t})^{T}}{\|{\bf{b}}_{n}-\mathbf{t}\|},\vspace{-1.5mm}$ (19)
so the scaled Jacobian is $\mathbf{J}(\mathbf{t}^{\prime})=\mathbf{J}(\mathbf{t})$. The original inverse-covariance $\mathbf{R}^{-1}$ is diagonal with entries of the form $1/(\|{\bf{b}}_{n}-\mathbf{t}\|^{\beta}\|{\bf{b}}_{m}-\mathbf{t}\|^{\beta})$ for the bi-static link from BS $n$ to a target and then to BS $m$. After scaling, we have:
$\|{\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime}\|=\kappa\,\|{\bf{b}}_{n}-\mathbf{t}\|\;\Longrightarrow\;\mathbf{R}^{\prime-1}=\kappa^{-2\beta}\,\mathbf{R}^{-1}.\vspace{-1.5mm}$ (20)
@@
$\mathbf{F}(\mathbf{t}^{\prime})=\mathbf{J}^{T}\bigl{(}\kappa^{-2\beta}\mathbf{R}^{-1}\bigr{)}\mathbf{J}=\kappa^{-2\beta}\,\mathbf{F}(\mathbf{t}).\vspace{-1.5mm}$ (21)
The CRLB is $\mathrm{trace}(\mathbf{F}^{-1})$. Thus, we have:
$\overline{\mathrm{CRLB}}\bigl{(}\mathcal{A}^{\prime},\{\mathbf{b}_{n}^{*}\}\bigr{)}=\kappa^{2\beta}\overline{\mathrm{CRLB}}\bigl{(}\mathcal{A},\{\mathbf{b}_{n}^{*}\}\bigr{)}.\vspace{-1.5mm}$ (22)
@@
Based on the conclusions of Propositions 1–4 above, we can further characterize the optimal deployment features for large-scale regions. In particular, for extensive areas the optimal deployment exhibits clear regularity (defined here as the presence of repetitive, periodic, or structured spatial arrangements), thereby significantly simplifying the design process.
@@
Area Periodicity. Although a rigorous proof in the general case is still open, it is intuitively reasonable that the optimal deployment $\mathbf{b}^{*}$ is $\Lambda$-periodic, whenever the sensing region is sufficiently large and BSs are densely deployed. In other words, for any translation vector $v\in\Lambda$, the optimal deployment follows the invariance to displacement principle as described in Proposition 1. The key intuition is that the total cost decomposes into a sum of identical local costs over each periodic sub-region, each depending mainly on the pattern within that sub-region and invariant under $\Lambda$-translations. If two sub-regions had different deployment patterns, one with higher cost, one could copy the lower-cost pattern into the other sub-regions to reduce the overall CRLB, contradicting the assumed optimality of $\mathbf{b}^{*}$. Therefore, the optimal deployment must repeat the same pattern across all sub-regions, yielding $\Lambda$-periodicity.
@@
Let $n_{0}$ be a fixed constant number of BSs optimally deployed in the region $\mathcal{A}\subset\mathbb{R}^{d}$, achieving optimal $\overline{\rm{CRLB}}=E^{*}(n_{0},\mathcal{A})$. If the BS count is increased by a factor of $N$, i.e., the total BS number becomes $N\times n_{0}$, then the optimal $\overline{\rm{CRLB}}$ satisfies a scaling law upper bound:
$\lim\limits_{N\to\infty}\frac{E^{*}(N\times n_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\,\geq\mathcal{O}\bigl{(}N^{-4/d}\bigr{)}.\vspace{-1.5mm}$ (23)
@@
Based on the analysis in Remark 2, in each subregion $\mathcal{A}_{i}$, we deploy $n_{0}$ BSs according to the optimal deployment strategy appropriately scaled to fit $\mathcal{A}_{i}$. By Proposition 4, scaling the region by a factor $\kappa$ results in the CRLB scaling by $\kappa^{2\beta}$. Hence each new subregion’s optimal CRLB is
$E^{*}(N\times n_{0},\mathcal{A})\leq$ $E^{*}(n_{0},\mathcal{A}_{i})$ (24)
$=$ $\kappa^{2\beta}\,E^{*}(n_{0},\mathcal{A})=N^{-4/d}\,E^{*}(n_{0},\mathcal{A}).$
The inequality in (24) is due to the fact that in practice all $N\times n_{0}$ BSs can cooperate fully. Hence this partition-based result is indeed an upper bound, and the actual cooperative CRLB can be further reduced. According to (24), the mean CRLB over $(N\times n_{0})$ BSs satisfies $E^{*}(Nn_{0},\mathcal{A})\leq N^{-4/d}\,E^{*}(n_{0},\mathcal{A})$, i.e.,
$\lim\limits_{N\to\infty}\frac{E^{*}(Nn_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\geq\mathcal{O}\bigl{(}N^{-4/d}\bigr{)}.\vspace{-1.5mm}$ (25)
@@
According to Theorem 1, the exponent $d=1,2,3$, yields the following explicit scaling law:
$\lim\limits_{N\to\infty}\frac{E^{*}(Nn_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\geq\begin{cases}\mathcal{O}(N^{-4}),&d=1,\\ \mathcal{O}(N^{-2}),&d=2,\\ \mathcal{O}(N^{-4/3}),&d=3.\end{cases}\vspace{-1.5mm}$ (26)
As the sensing-region dimensionality $d$ increases, nodes become dispersed across more axes, thereby reducing the relative performance improvement contributed by each BS.
@@
The above $\mathcal{O}(N^{-2\beta/d})$ bound in Theorem 1 is relatively tight under dense BS environments: As $N$ grows, subregions more distant from a given target contribute vanishingly small error due to the increased distance aggravating the path-loss. In particular, when $n_{0}$ is large so that each original BS region is already small, further subdividing introduces only marginal gains from distant cells, implying that the actual cooperative CRLB closely approaches this upper bound. Therefore, it follows that
$\lim\limits_{n_{0}\to\infty}\lim\limits_{N\to\infty}\frac{E^{*}(N\times n_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\,\approx\mathcal{O}\bigl{(}N^{-2\beta/d}\bigr{)}\vspace{-1.5mm}$ (27)
@@ III-C Low-complexity Sensing BS Deployment Optimization
Based on the Propositions and Theorem above, we can determine the optimal deployment plan for a new region by constructing a database of optimal configurations. By comparing the shape of the new region to those stored in the database, and using displacement, rotation, and scaling adjustments as necessary, one can readily construct deployment solutions. Specifically, for any given deployment region $\mathcal{A}$, there exists a set of representative regions $\{\mathcal{A}_{i}\}_{i\in\cal{{I}}}$ for which the optimal deployments have already been solved, along with the corresponding solution set $\{{\bf{b}}_{n}^{*}\}_{i\in\cal{{I}}}$. Then, there exists a transformation consisting of displacement, rotation, and scaling that maps $\mathcal{A}_{i}$ to $\mathcal{A}$, where $i$ denotes the index of the candidate area set $\cal{{I}}$. Then, the corresponding optimal deployment ${\bf{b}}^{*}$ can be derived using the inverse of the transformations applied above:
${\bf{b}}^{*}=\kappa\mathbf{O}_{H}\mathbf{R}_{\varphi}{\bf{b}}_{i}^{*}+\mathbf{\Delta}.\vspace{-1.5mm}$ (28)
@@ IV-A Problem Reformulation
Directly integrating localization performance over a continuous region requires averaging over an uncountable set of target positions and random channel realizations, and this thus yields high-dimensional integrals that admit no closed-form solution. To handle the continuous-region integration, we acquire $K$ sampled points $\{\boldsymbol{t}_{k}\}_{k=1}^{K}$ uniformly distributed over the area of interest. Similarly, as for the communication constraint, the communication performance is evaluated at the sampled points drawn, ensuring that the area communication capacity meets the required threshold $R^{th}$.
Because the instantaneous channel gains $g_{n}$ in (10) fluctuate randomly due to small-scale fading, it is impractical to adjust the BS placement relying on these transient values. To address the infeasibility of optimizing over random channel realizations, we first attain the average beamforming gain for both communication and sensing by taking the expectation over the small-scale fading, thereby ensuring a robust, long-term gain in both communication and sensing. Then, we can derive the distribution of $g_{j}$ based on the moment matching technique of [34]. In (10), $g_{n}=p^{c}\left|\mathbf{h}_{n}^{H}\mathbf{w}_{n}^{c}\right|^{2}$ represents the effective desired signals’ channel gain, where $g_{n}\sim\Gamma\left(M_{\mathrm{t}}-1,p^{c}\right)$ [34, 18]. Then, we have ${\rm{E}}[\left|\mathbf{h}_{j}^{H}\mathbf{w}_{j}^{c}\right|^{2}]=p^{c}$. Similarly, the expected transmit beamforming gain for sensing can be formulated as ${\rm{E}}[\left|\mathbf{a}^{H}(\theta_{n})\mathbf{w}_{n}^{s}\right|^{2}]=M_{\rm{t}}-1$, i.e., ${\rm{E}}[G_{t}]=M_{\rm{t}}-1$.
@@
This sampling-based reformulation transforms the original continuous optimization problem into a tractable discrete form. Moreover, the propositions and theorem presented in Section III remain applicable in the discrete case.
@@ IV-B Proposed Problem Solution
Jointly optimizing multiple BSs’ location poses significant challenges, primarily due to the highly non-convex nature of the CRLB objective function. To tackle this challenge, we adopt the MM algorithmic framework to facilitate problem solving. This iterative optimization method proves effective in handling non-convex optimization problems by sequentially majorizing the objective function $f(x)$ with a surrogate function $g(x\mid x_{t})$.
@@ IV-B2 Problem Transformation
Consider the optimization problem, where we aim for improving the localization performance, starting with an initialization that ensures having no node lies in the same plane as the target point. This initialization condition is crucial for achieving enhanced localization accuracy. In the following, we present a deployment optimization problem transformation method based on the MM framework. At each iteration, we optimize the position of one BS, while holding the other $(N-1)$ BSs fixed, thereby decomposing the original high-dimensional problem into a sequence of simpler subproblems.
In the following, we introduce an optimization algorithm designed for updating the location of BS $n$ as a representative example. For notational simplicity, we first reformulate the problem for a single target point. The extension to the general case having $K$ target points can be achieved by summing the resultant objective function over all targets. To separate the term related to BS $n$ in the objective function, we define ${{\mathbf{v}}_{n,k}}=\frac{{{{{{\bf{b}}_{n}}-{\mathbf{t}_{k}}}}}}{{\left\|{{{{\bf{b}}_{n}}}-{\mathbf{t}_{k}}}\right\|}}$. Then the CRLB of target point ${\bf{t}}_{k}$, under centralized fusion-based localization, can be formulated as
${\mathrm{CRLB}}({\bf{t}}_{k})=\text{trace}\left(\mathbf{F}\left({\bf{t}}_{k}\right)^{-1}\right)$ (31)
@@
where ${\bf{M}}=\sum_{\begin{subarray}{c}j=1,j\neq n\end{subarray}}^{N}\sum_{\begin{subarray}{c}i=1,i\neq n\end{subarray}}^{N}\frac{({\mathbf{v}}_{i,k}+{\mathbf{v}}_{j,k})({\mathbf{v}}_{i,k}^{T}+{\mathbf{v}}_{j,k}^{T})}{\|{{\bf{b}}}_{i}-{\mathbf{t}}_{k}\|^{\beta}\|{{\bf{b}}}_{j}-{\mathbf{t}}_{k}\|^{\beta}}$, including the term unrelated to the location of BS $n$.
To facilitate problem transformation, let ${\bf{D}}_{n,k}=\mathrm{diag}\!\left(\!\frac{2}{{\left\|{\bf{b}}_{1}-{\bf{t}}_{k}\right\|^{\beta}\left\|{\bf{b}}_{n}-{\bf{t}}_{k}\right\|^{\beta}}},\dots,\frac{1}{{\left\|{\bf{b}}_{n}-{\bf{t}}_{k}\right\|^{2\beta}}},\dots,\frac{2}{{\left\|{\bf{b}}_{N}-{\bf{t}}_{k}\right\|^{\beta}\!\left\|{\bf{b}}_{n}-{\bf{t}}_{k}\right\|^{\beta}}}\!\!\right)$ and ${\bf{U}}_{n,k}=\left[{\mathbf{v}}_{1,k}+{\mathbf{v}}_{n,k},\dots\mathbf{v}_{n,k}+{\mathbf{v}}_{n,k},\dots\mathbf{v}_{N,k}+{\mathbf{v}}_{n,k}\right]$. Then we can reformulate the sensing link terms in (31) related to BS $n$ into ${\bf{U}}_{n,k}{\bf{D}}_{n,k}{{\bf{U}}}_{n,k}^{T}=\frac{4{\mathbf{v}}_{n,k}{\mathbf{v}}_{n,k}^{T}}{\|{{\bf{b}}}_{n}-{\mathbf{t}}_{k}\|^{2\beta}}+2\sum_{\begin{subarray}{c}i=1,i\neq n\end{subarray}}^{N}\frac{({\mathbf{v}}_{i,k}+{\mathbf{v}}_{n,k})({\mathbf{v}}_{i,k}^{T}+{\mathbf{v}}_{n,k}^{T})}{\|{{\bf{b}}}_{i}-{\mathbf{t}}_{k}\|^{\beta}\|{{\bf{b}}}_{n}-{\mathbf{t}}_{k}\|^{\beta}}$. According to the Schur complement [35], it follows that
@@
$=$ ${\mathrm{Tr}}\left({\bf{M}}^{-1}-{\bf{M}}^{-1}{\bf{U}}_{n,k}{\bf{Z}}_{n,k}^{-1}{\bf{U}}_{n,k}^{T}{\bf{M}}^{-1}\!\right),\vspace{-1.5mm}$
where ${\bf{Z}}_{n,k}={\bf{D}}_{n,k}^{-1}+{\bf{U}}_{n,k}^{T}{\bf{M}}^{-1}{\bf{U}}_{n,k}$. Then, the problem can be equivalently reformulated as ${\mathrm{Tr}}\left({\bf{M}}^{-1}{\bf{U}}_{n,k}{\bf{Z}}_{n,k}^{-1}{\bf{U}}_{n,k}^{T}{\bf{M}}^{-1}\right)$. To further facilitate the problem solution, the following inequality is introduced:
@@
where $\mathbf{G}_{n,k}=\left((\mathbf{Z}_{n,k}^{(r)})^{-1}(\mathbf{U}_{n,k}^{(r)})^{T}{\mathbf{M}}^{-2}\mathbf{U}_{n,k}^{(r)}(\mathbf{Z}_{n,k}^{(r)})^{-1}\right)^{T}$ and $\mathbf{P}_{n,k}=2{\mathbf{M}}^{-2}\mathbf{U}_{n,k}^{(r)}(\mathbf{Z}_{n,k}^{(r)})^{-1}$. Then, by substituting ${\bf{Z}}_{n,k}$ into (34), the objective function in (34) can be expressed as $\operatorname{Tr}\left(\mathbf{G}_{n,k}{\mathbf{D}}_{n,k}^{-1}\right)+\operatorname{Tr}\left(\mathbf{G}_{n,k}\mathbf{U}_{n,k}^{T}{\mathbf{M}}^{-1}\mathbf{U}_{n,k}\right)-\operatorname{Tr}\left(\mathbf{P}_{n,k}^{T}\mathbf{U}_{n,k}\right)$, where $\operatorname{Tr}\left(\mathbf{G}_{n,k}{\mathbf{D}}_{n,k}^{-1}\right)$ comprises quadratic and quartic functions of the BS location vectors. By introducing an appropriate change of variables $\delta_{n,k}$ satisfying $\left\|\mathbf{t}_{k}-{\bf{b}}_{n}\right\|^{\beta}\leq\delta_{n,k},\forall n,k$, it can be rendered convex. Then, we can first handle the term $\operatorname{Tr}\left(\mathbf{G}_{n,k}\mathbf{U}_{n,k}^{T}{\mathbf{M}}^{-1}\mathbf{U}_{n,k}\right)$ as follows. Define $\mathbf{U}_{n,k}^{c}=\left[\mathbf{v}_{1,k},\dots,\mathbf{v}_{n-1,k},0,\mathbf{v}_{n+1,k},\dots,\mathbf{v}_{N,k}\right]^{T}$. Then, by substituting ${\bf{Z}}_{n,k}$ into the above equation and separating the linear and quadratic contributions in $\mathbf{v}_{n,k}$, the objective function can be transformed into

### v2

@@ III-B Scaling Law Analysis
$\frac{({\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime})^{T}}{\|{\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime}\|}=\frac{(\kappa{\bf{b}}_{n}-\kappa\mathbf{t})^{T}}{\kappa\|{\bf{b}}_{n}-\mathbf{t}\|}=\frac{({\bf{b}}_{n}-\mathbf{t})^{T}}{\|{\bf{b}}_{n}-\mathbf{t}\|},\vskip 0.0pt$ (19)
so the scaled Jacobian is $\mathbf{J}(\mathbf{t}^{\prime})=\mathbf{J}(\mathbf{t})$. Ignoring fixed multiplicative constants, the original inverse‐covariance $\mathbf{R}^{-1}$ is diagonal with entries of the form $1/(\|{\bf{b}}_{n}-\mathbf{t}\|^{\beta}\|{\bf{b}}_{m}-\mathbf{t}\|^{\beta})$ for the bi-static link from BS $n$ to a target and then to BS $m$. After scaling, we have:
$\|{\bf{b}}_{n}^{\prime}-\mathbf{t}^{\prime}\|=\kappa\,\|{\bf{b}}_{n}-\mathbf{t}\|\;\Longrightarrow\;\mathbf{R}^{\prime-1}=\kappa^{-2\beta}\,\mathbf{R}^{-1}.\vskip 0.0pt$ (20)
@@
$\mathbf{F}(\mathbf{t}^{\prime})=\mathbf{J}^{T}\bigl(\kappa^{-2\beta}\mathbf{R}^{-1}\bigr)\mathbf{J}=\kappa^{-2\beta}\,\mathbf{F}(\mathbf{t}).\vskip 0.0pt$ (21)
The CRLB is $\mathrm{Tr}(\mathbf{F}^{-1})$. Thus, we have:
$\overline{\mathrm{CRLB}}\bigl(\mathcal{A}^{\prime},\{\mathbf{b}_{n}^{*}\}\bigr)=\kappa^{2\beta}\overline{\mathrm{CRLB}}\bigl(\mathcal{A},\{\mathbf{b}_{n}^{*}\}\bigr).\vskip 0.0pt$ (22)
@@
By exploiting invariance to rigid motions (global translations, rotations, and reflections) together with the uniform-area scaling in Proposition 4, an optimal BS deployment computed for a canonical region can transfer directly to any translated/rotated/reflected/scaled instance of that region. Based on the conclusions of Propositions 1-4 above, we can further characterize the optimal deployment features for large-scale regions. In particular, for extensive areas the optimal deployment exhibits clear regularity (defined here as the presence of repetitive, periodic, or structured spatial arrangements), thereby significantly simplifying the design process.
@@
Area Periodicity. Although a rigorous proof in the general case is still open, it is intuitively reasonable that the optimal deployment $\mathbf{b}^{*}$ is $\Lambda$-periodic, whenever the sensing region is sufficiently large, effectively a tiling of identical subregions, and BSs are densely deployed. In other words, for any translation vector $v\in\Lambda$, the optimal deployment follows the invariance to displacement principle as described in Proposition 1. The key intuition is that the total cost decomposes into a sum of identical local costs over each periodic sub-region, each depending mainly on the pattern within that sub-region and invariant under $\Lambda$-translations. If two sub-regions had different deployment patterns, one with higher cost, one could copy the lower-cost pattern into the other sub-regions to reduce the overall CRLB, contradicting the assumed optimality of $\mathbf{b}^{*}$. Therefore, the optimal deployment must repeat the same pattern across all sub-regions, yielding $\Lambda$-periodicity.
@@
Let $n_{0}$ be a fixed constant number of BSs optimally deployed in a sufficiently large region $\mathcal{A}\subset\mathbb{R}^{d}$, achieving optimal $\overline{\rm{CRLB}}=E^{*}(n_{0},\mathcal{A})$. If the BS count is increased by a factor of $N$, i.e., the total BS number becomes $N\times n_{0}$, then the optimal $\overline{\rm{CRLB}}$ satisfies a scaling law upper bound:
$\lim\limits_{N\to\infty}\frac{E^{*}(N\times n_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\,\leq\mathcal{O}\bigl(N^{-2\beta/d}\bigr).\vskip 0.0pt$ (23)
@@
Let $E^{*}(n,\mathcal{A})$ denote the minimum achievable area-CRLB for the sensing-only deployment problem (15) when $n$ BSs are deployed over region $\mathcal{A}$. Based on Remark 2, we partition $\mathcal{A}$ into $N$ congruent subregions $\{\mathcal{A}_{i}\}_{i=1}^{N}$, and in each $\mathcal{A}_{i}$ we place $n_{0}$ BSs by taking the optimal deployment for $\mathcal{A}$ and uniformly scaling it to fit $\mathcal{A}_{i}$. This construction yields a feasible deployment over the entire region $\mathcal{A}$ with a corresponding achieved objective value, denoted by ${E}_{\rm cons}(N\times n_{0},\mathcal{A})$ for each subregion. By Proposition 4, scaling the region by a factor $\kappa$ results in the CRLB scaling by $\kappa^{2\beta}$. Hence each new subregion’s optimal CRLB is
$E^{*}(N\times n_{0},\mathcal{A})\leq$ ${E}_{\rm cons}(N\times n_{0},\mathcal{A})$ (24)
$\overset{({\rm{a}})}{\leq}$ $E^{*}(n_{0},\mathcal{A}_{i})$
$=$ $\kappa^{2\beta}\,E^{*}(n_{0},\mathcal{A})=N^{-2\beta/d}\,E^{*}(n_{0},\mathcal{A}).$
The inequality $({\rm a})$ holds due to the fact that in practice all $N\times n_{0}$ BSs can cooperate fully. Hence this partition-based result is indeed an upper bound, and the actual cooperative CRLB can be further reduced. According to (24), the mean CRLB over $(N\times n_{0})$ BSs satisfies $E^{*}(Nn_{0},\mathcal{A})\leq N^{-2\beta/d}\,E^{*}(n_{0},\mathcal{A})$, i.e.,
$\lim\limits_{N\to\infty}\frac{E^{*}(Nn_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\leq\mathcal{O}\bigl(N^{-2\beta/d}\bigr).\vskip 0.0pt$ (25)
@@
According to Theorem 1, when $\beta=2$, the exponent $d=1,2,3$, yields the following explicit scaling law:
$\lim\limits_{N\to\infty}\frac{E^{*}(Nn_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\leq\begin{cases}\mathcal{O}(N^{-4}),&d=1,\\ \mathcal{O}(N^{-2}),&d=2,\\ \mathcal{O}(N^{-4/3}),&d=3.\end{cases}\vskip 0.0pt$ (26)
With BS count/density fixed, increasing the sensing-region dimension $d$ spreads nodes across more axes, shrinking the average solid-angle separation between BS look orientation and lengthening typical BS-target ranges; this dilutes bearing diversity, weakens the Fisher information, and reduces the marginal A-CRLB improvement per added BS.
@@
The above $\mathcal{O}(N^{-2\beta/d})$ bound in Theorem 1 is relatively tight under dense BS environments: As $N$ grows, subregions more distant from a given target contribute vanishingly small error due to the increased distance aggravating the path‐loss. In particular, when $n_{0}$ is large so that each original BS region is already small, further subdividing introduces only marginal gains from distant cells, implying that the actual cooperative CRLB closely approaches this upper bound. Therefore, it follows that
$\lim\limits_{n_{0}\to\infty}\lim\limits_{N\to\infty}\frac{E^{*}(N\times n_{0},\mathcal{A})}{E^{*}(n_{0},\mathcal{A})}\,\approx\mathcal{O}\bigl(N^{-2\beta/d}\bigr)\vskip 0.0pt$ (27)
@@ III-C Low-complexity Sensing BS Deployment Optimization
[…]
${\bf{b}}^{*}=\kappa\mathbf{O}_{H}\mathbf{R}_{\varphi}{\bf{b}}_{i}^{*}+\mathbf{\Delta}.\vskip 0.0pt$ (28)
@@ IV-A Problem Reformulation
Directly integrating localization performance over a continuous region requires averaging over an uncountable set of target positions and random channel realizations, and this thus yields high-dimensional integrals that admit no closed-form solution. To handle the continuous‐region integration, we acquire $K$ sampled points $\{\boldsymbol{t}_{k}\}_{k=1}^{K}$ uniformly distributed over the area of interest. Similarly, as for the communication constraint, the communication performance is evaluated at the sampled points drawn, ensuring that the area communication capacity meets the required threshold $R^{th}$.
Because the instantaneous channel gains $g_{n}$ in (10) fluctuate randomly due to small-scale fading, it is impractical to adjust the BS placement relying on these transient values. To address the infeasibility of optimizing over random channel realizations, we first attain the average beamforming gain for both communication and sensing by taking the expectation over the small-scale fading, thereby ensuring a robust, long-term gain in both communication and sensing. Then, we can derive the distribution of $g_{j}$ based on the moment matching technique of [37]. In (10), $g_{n}=\left|\mathbf{h}_{n}^{H}\mathbf{w}_{n}^{c}\right|^{2}$ represents the effective desired signals’ channel gain, where $g_{n}\sim\Gamma\left(M_{\mathrm{t}}-1,p^{c}\right)$ [37, 18]. Similarly, the expected transmit beamforming gain for sensing can be formulated as ${\rm{E}}[\left|\mathbf{a}^{H}_{M_{\rm{t}}}(\Omega_{n})\mathbf{w}_{n}^{s}\right|^{2}]=p^{s}\left(M_{\rm{t}}-1\right)$, i.e., ${\rm{E}}[G_{t}]=p^{s}\left(M_{\rm{t}}-1\right)$.
@@
This sampling-based reformulation transforms the original continuous optimization problem into a tractable discrete form. The invariances in the preceding propositions and theorem carry over to the discretized formulation: co-transforming the sampling set with the region preserves sample-BS geometry, leaving the discrete objective and constraints unchanged. Hence the results of Section III apply verbatim.
@@ IV-B Proposed Problem Solution
Jointly optimizing multiple BS locations is challenging because the CRLB objective is highly non-convex. Plain gradient descent algorithm, without a majorizing surrogate, is markedly step-size-sensitive and often fails to achieve monotone descent. To tackle this challenge, we adopt a block-coordinate (alternating) MM scheme: at iteration $t$, with N1N-1 BSs fixed, we majorize $f(x)$ by a tight surrogate function $g(x\mid x_{t})$ and minimize it using a short projected gradient-descent step. This yields monotonic descent and convergence to a stationary point under standard MM conditions.
@@ IV-B2 Problem Transformation
Consider the optimization problem, where we aim for improving the localization performance, starting with an initialization that ensures having no node lies in the same plane as the target point. This initialization condition is crucial for achieving enhanced localization accuracy. In the following, we present a deployment optimization problem transformation method based on the MM framework. At each iteration, we optimize the position of one BS, while holding the other (N1)(N-1) BSs fixed, thereby decomposing the original high-dimensional problem into a sequence of simpler subproblems.
In the following, we introduce an optimization algorithm designed for updating the location of BS $n$ as a representative example. For notational simplicity, we first reformulate the problem for a single target point. The extension to the general case having $K$ target points can be achieved by summing the resultant objective function over all targets. To separate the term related to BS $n$ in the objective function, we define ${{\mathbf{v}}_{n,k}}=\frac{{{{{{\bf{t}}_{k}}-{\mathbf{b}_{n}}}}}}{{\left\|{{{{\bf{t}}_{k}}}-{\mathbf{b}_{n}}}\right\|}}$. Then the CRLB of target point ${\bf{t}}_{k}$, under centralized fusion-based localization, can be formulated as
${\mathrm{CRLB}}({\bf{t}}_{k})=\text{Tr}\left(\mathbf{F}\left({\bf{t}}_{k}\right)^{-1}\right)$ (31)
@@
where ${\bf{M}}=\sum_{\begin{subarray}{c}j=1,j\neq n\end{subarray}}^{N}\sum_{\begin{subarray}{c}i=1,i\neq n\end{subarray}}^{N}\frac{({\mathbf{v}}_{i,k}+{\mathbf{v}}_{j,k})({\mathbf{v}}_{i,k}^{T}+{\mathbf{v}}_{j,k}^{T})}{\|{{\bf{b}}}_{i}-{\mathbf{t}}_{k}\|^{\beta}\|{{\bf{b}}}_{j}-{\mathbf{t}}_{k}\|^{\beta}}$, including the term unrelated to the location of BS $n$. Assume that the initialization yields a non-degenerate BS-target geometry, so that the resulting matrix $\mathbf{M}$ is positive definite and hence invertible.
To facilitate problem transformation, let ${\bf{D}}_{n,k}=\mathrm{diag}\!\left(\!\frac{2}{{\left\|{\bf{b}}_{1}-{\bf{t}}_{k}\right\|^{\beta}\left\|{\bf{b}}_{n}-{\bf{t}}_{k}\right\|^{\beta}}},\dots,\frac{1}{{\left\|{\bf{b}}_{n}-{\bf{t}}_{k}\right\|^{2\beta}}},\dots,\frac{2}{{\left\|{\bf{b}}_{N}-{\bf{t}}_{k}\right\|^{\beta}\!\left\|{\bf{b}}_{n}-{\bf{t}}_{k}\right\|^{\beta}}}\!\!\right)$ $\in\mathbb{R}^{N\times N}$ and ${\bf{U}}_{n,k}=\left[{\mathbf{v}}_{1,k}+{\mathbf{v}}_{n,k},\dots,{\mathbf{v}}_{N,k}+{\mathbf{v}}_{n,k}\right]\in\mathbb{R}^{3\times N}$. Then we can reformulate the sensing link terms in (31) related to BS $n$ into ${\bf{U}}_{n,k}{\bf{D}}_{n,k}{{\bf{U}}}_{n,k}^{T}=\frac{4{\mathbf{v}}_{n,k}{\mathbf{v}}_{n,k}^{T}}{\|{{\bf{b}}}_{n}-{\mathbf{t}}_{k}\|^{2\beta}}+2\sum_{\begin{subarray}{c}i=1,i\neq n\end{subarray}}^{N}\frac{({\mathbf{v}}_{i,k}+{\mathbf{v}}_{n,k})({\mathbf{v}}_{i,k}^{T}+{\mathbf{v}}_{n,k}^{T})}{\|{{\bf{b}}}_{i}-{\mathbf{t}}_{k}\|^{\beta}\|{{\bf{b}}}_{n}-{\mathbf{t}}_{k}\|^{\beta}}$. According to the Woodbury identity [38], it follows that
@@
$=$ ${\mathrm{Tr}}\left({\bf{M}}^{-1}-{\bf{M}}^{-1}{\bf{U}}_{n,k}{\bf{Z}}_{n,k}^{-1}{\bf{U}}_{n,k}^{T}{\bf{M}}^{-1}\!\right),\vskip 0.0pt$
where ${\bf{Z}}_{n,k}={\bf{D}}_{n,k}^{-1}+{\bf{U}}_{n,k}^{T}{\bf{M}}^{-1}{\bf{U}}_{n,k}$. Then, the problem can be equivalently reformulated as maximizing ${\mathrm{Tr}}\left({\bf{M}}^{-1}{\bf{U}}_{n,k}{\bf{Z}}_{n,k}^{-1}{\bf{U}}_{n,k}^{T}{\bf{M}}^{-1}\right)$. To further facilitate the problem solution, the following inequality is introduced:
@@
where $\mathbf{G}_{n,k}=\left((\mathbf{Z}_{n,k}^{(r)})^{-1}(\mathbf{U}_{n,k}^{(r)})^{T}{\mathbf{M}}^{-2}\mathbf{U}_{n,k}^{(r)}(\mathbf{Z}_{n,k}^{(r)})^{-1}\right)^{T}$ and $\mathbf{P}_{n,k}=2{\mathbf{M}}^{-2}\mathbf{U}_{n,k}^{(r)}(\mathbf{Z}_{n,k}^{(r)})^{-1}$. Then, by substituting ${\bf{Z}}_{n,k}$ into (34), the objective function in (34) can be expressed as $\operatorname{Tr}\left(\mathbf{G}_{n,k}{\mathbf{D}}_{n,k}^{-1}\right)+\operatorname{Tr}\left(\mathbf{G}_{n,k}\mathbf{U}_{n,k}^{T}{\mathbf{M}}^{-1}\mathbf{U}_{n,k}\right)-\operatorname{Tr}\left(\mathbf{P}_{n,k}^{T}\mathbf{U}_{n,k}\right)$, where $\operatorname{Tr}\left(\mathbf{G}_{n,k}{\mathbf{D}}_{n,k}^{-1}\right)$ comprises quadratic and quartic functions of the BS location vectors. By introducing an appropriate change of variables $\delta_{n,k}$ satisfying $\left\|\mathbf{t}_{k}-{\bf{b}}_{n}\right\|^{\beta}\leq\delta_{n,k},\forall n,k$, it can be rendered convex. Then, we can first handle the term $\operatorname{Tr}\left(\mathbf{G}_{n,k}\mathbf{U}_{n,k}^{T}{\mathbf{M}}^{-1}\mathbf{U}_{n,k}\right)$ as follows. Define $\mathbf{U}_{n,k}^{c}=\left[\mathbf{v}_{1,k},\dots,\mathbf{v}_{n-1,k},0,\mathbf{v}_{n+1,k},\dots,\mathbf{v}_{N,k}\right]$, we have $\mathbf{U}_{n,k}=\mathbf{U}_{n,k}^{c}+\mathbf{v}_{n,k}\mathbf{1}^{T}+\mathbf{v}_{n,k}\mathbf{e}_{n}^{T}$. Then, by substituting ${\bf{Z}}_{n,k}$ into the above equation and separating the linear and quadratic contributions in $\mathbf{v}_{n,k}$, the objective function can be transformed into